A coil with resistance ‘

6

Ω

,

and

inductance

8

’ mH

is connected

in parallel

with a

series

combination of

a

resist

or and

capacitor

having

resistance R

Ω and

reactance X

c

Ω.

If

t

he equivalent

impedance

of the resulting circuit is

Z =

5

/0

o

Ω at

ω

=1000 rad/sec

. Find the

values of R & C

.



1
Expert's answer
2021-03-11T06:56:14-0500

Given, "Z_L = 6+j8"


It is connected across "R-jX"


The resultant parallel combination yields

"5\angle0" .


i. e"\frac{(6+j8) (R-jX) }{(6+R) +j(8-X)} =5+j0"


Simplifying and equating real and imaginary parts, we have,


"R+8X=30 \newline
8R-X=40"


Solving we have

"R= 5.38\Omega\newline X=3.077\Omega \newline 
Also Z = 5.38-j3.07"


Also "X=1/C\omega\implies C=0.325\mu F"